Re:No person who was born blind has ever been diagnosed with...
by primo4444 on Wednesday February 12, 2020 @08:49AM
(#22308759)
No, the .02 is correct. If you ignore the percentages part (since both left and right side use it), you're doing
If you do it as probabilities not expressed as percentages, it's 0.0072 x 0.0003, which is 0.0002, but that's 0.02%
[ Parent
| on HN ]
Re:No person who was born blind has ever been diagnosed with...
by rlpb on Wednesday February 12, 2020 @08:57AM
(#22308814)
You can't just ignore the percentage signs.
0.72% expressed as a decimal is 0.0072. 0.03% expressed as a decimal is 0.0003. 0.0072 x 0.0003 = 0.000002. Expressed back as a percentage, that's 0.0002%.
Re:No person who was born blind has ever been diagnosed with...
by gus_massa on Wednesday February 12, 2020 @08:58AM
(#22308820)
> it's 0.0072 x 0.0003, which is 0.0002, but that's 0.02%
No. 0.0072 x 0.0003 = 0.00000216 ~= 0.000002 = 0.0002%
Re:No person who was born blind has ever been diagnosed with...
by primo4444 on Wednesday February 12, 2020 @09:02AM
(#22308847)
You're right. Not sure how I fell into that trap so easily.
Re:No person who was born blind has ever been diagnosed with...
by injb on Wednesday February 12, 2020 @09:49AM
(#22309193)
I wonder if it's better to treat percentage as a unit or variable. So a-percent x b-percent = c(percent-squared), or c divided by 100 twice.
Re:No person who was born blind has ever been diagnosed with...
by petrogradphilos on Wednesday February 12, 2020 @11:11AM
(#22309992)
You could also replace % with 10⁻² and use scientific notation:
0.72% ⋅ 0.03%
= 0.72 ⋅ 10⁻² ⋅ 0.03 ⋅ 10⁻²
= 7.2 ⋅ 10⁻³ ⋅ 3 ⋅ 10⁻⁴
= 7.2 ⋅ 3 ⋅ 10⁻³ ⋅ 10⁻⁴
= 21.6 ⋅ 10⁻⁷
= 2.16 ⋅ 10⁻⁶
≈ 2 in 1 million